OpenStudy (anonymous):

A 1266.9 kg car is traveling at 25.3 m/s when the driver takes his foot off the gas pedal. It takes 4.6 s for the car to slow down to 20 m/s. How large is the net force slowing the car?

OpenStudy (anonymous):

According to newton's second law, F-r=ma, where F is driving force, R is the resistive force , m is mass and a is acceleration. According to your question, there is no driving froce. Only R acts to slow the car down. Plus, you should know that a is the same as rate of change of velocity or (v-u)/t. Replacing the values that you have= a= (20-25.3)/4.6=-1.15m/s^2. Going back to F-R=ma, becomes now 0-R=1266.9*(-1.15) and R becomes 1456.9N.